Example 4: Mixing Cleaning Solutions.
King’s Service Station uses two kinds of cleaning solution containing acid and water. “Attack” is 2% acid and “Blast” is 6% acid. They want to mix the two to get 60 qt of a solution that is 5% acid. How many quarts of each should they use?
Thank you!!!

Answer by ptaylor(2198)

Let x=amount of 2% acid needed
Then 60-x=amount of 6% acid needed
Now we know that the total amount of pure acid before the mixture ((0.02x+0.06(60-x))has to equal the total amount of pure acid after the mixture (0.05*60). So our equation to solve is:
0.02x+0.06(60-x)=0.05*60 simplify
0.02x+3.6-0.06x=3 subtract 3.6 from each side
-0.04x=-0.6 divide each side by -0.04
x=15 qts--------------------amt of 2% acid needed
60-x=60-15=45 qts--------------amt of 6% acid needed
CK
0.02*15+0.06*45=0.05*60
0.3+2.7=3
3=3
Hope this helps----ptaylor